# WaterPark: A Robustness Assessment of Language Model Watermarking

### v1

@@ Appendix A Parameter Setting
Appendix B Watermark Effectiveness
We first evaluate the effectiveness of different watermarkers. Following prior work[26, 49, 29], for each watermarker, we sample 1,000 prompts and use the LLM in combination with the watermarker to generate the watermarked texts; meanwhile, we select human responses to the same prompts as the non-watermarked texts. We then measure the accuracy of the watermarker’s detector in distinguishing the watermarked and non-watermarked texts.
B.1 Overall Effectiveness
@@
Meanwhile, most methods observe marginal performance drops on Llama2-7B, as shown in Figure 14 (c-d). For instance, compared with its performance on OPT-1.3B, the AUC of SIR drops by 0.11 and 0.08 on C4 and HC3 respectively. This observation aligns with previous research [29], indicating that watermarkers are more effective on OPT compared with Llama2. This phenomenon can be partly understood through the following explanation. In contrast of less capable LLMs (e.g., OPT-1.3B), Llama2-7B typically produces texts of lower perplexity. Since most watermarkers inject watermarks by slightly altering the next-token distribution, the lower perplexity in Llama2’s outputs hampers the effectiveness of such perturbations. Moreover, it is observed that the performance of various methods varies significantly across different datasets. For instance, UG’s AUC differs by 0.28 between C4 and HC3, while UB’s AUC differs by 0.11. This observation is further supported by the TPR measures at fixed FPRs (fixed as 1%), as shown in Figure 14 (e-f). Our findings suggest that the concrete dataset/task tends to have a larger impact on watermarkers over more capable LLMs.
Observation 1 – Ensuring watermarkers achieve the same quality and effectiveness as reported in their works and existing watermarkers tend to be more effective on LLMs with higher perplexity.
B.2 Impact of Text Length
@@
Observation 2 – Different watermarkers have varying sensitivity to text length; distribution transform-based watermarkers are the least sensitive.
B.3 Impact of Temperature
@@
Figure 16 compares how the TPRs of different watermarkers vary with the setting of $\tau=$ 0.7, 1.0, and 1.3 (with FPRs fixed as 1%). Observe that the performance of most watermarkers marginally improves with $\tau$, which corroborates prior work[49]. For instance, the TPR of TGRL increases by about 0.05 as $\tau$ varies from 0.7 to 1.3. Interestingly, in contrast, the TPR of UPV decreases as $\tau$ grows. This can be explained as follows. UPV employs a neural network as the detector that depends on general textual features, which tends to be more sensitive to increasing randomness, compared with other watermarkers that rely on specific watermark signals (e.g., green/red-listed tokens).
Observation 3 – The effectiveness of watermarkers that rely on explicit watermark signals improves with the temperature.
Appendix C Watermark Fidelity
@@
We evaluate the impact of different watermarkers on the text quality. We compare the original text $T$ and watermarked text $\tilde{T}$ using the metrics (detailed in §4) of BERTScore[75], P-SP[65], and MAUVE[50]. The results are summarized in Figure 17.
@@
Observation 4 – There exists a fundamental trade-off between the watermarker’s fidelity and indistinguishability.
Appendix D Additional Results
D.1 Multi-bit Watermarking
While our study focuses on one-bit watermarkers, for completeness, we also evaluate CTWL[60], a multi-bit watermarker. In contrast of one-bit watermarkers that encode only a single bit of information (i.e., whether a given text is watermarked), a multi-bit watermarker can encode multiple bits of information into the watermarked text, such as the generating model, the date of generation, and other details. However, despite its larger information capacity, we find that a multi-bit watermarker is typically less robust compared to one-bit watermarkers, as illustrated in Figure 18 and  19.
@@ D.1 Multi-bit Watermarking
D.2 Robustness of Watermarkers on Llama2
Our evaluation in §5 mainly uses OPT-1.3B as the underlying LLM. Here, we present the results on Llama2-7B, a more capable LLM.
D.2.1 Linguistic Variation Attack
@@
Figure 20: Resilience of LLM watermarkers on Llama2 against linguistic variation attacks.
D.2.2 Lexical Editing Attack
@@
Figure 21: Resilience of LLM watermarkers on Llama2 against lexical editing attacks.
D.2.3 Text-Mixing Attacks
@@
Figure 22: Resilience of LLM watermarkers on Llama2 against text-mixing attacks.
D.2.4 Paraphrasing Attack
@@
Figure 23: Resilience of LLM watermarkers on Llama2 against paraphrasing attacks.
D.3 Fidelity Preservation

### v2

@@ 6 Conclusion
Limitations
We now examine the limitations of our study and identify promising directions for future research.
Watermarkers. This study primarily focuses on training-free, pre-generation watermarking, which is applicable to any given LLM and offers flexible control over multiple criteria (e.g., quality, effectiveness, and robustness). While we have included peer-reviewed watermarking methods to the best of our ability, many methods, such as training-time watermarking and post-generation watermarking, are not covered in this paper.
Threat Model. Notably, our evaluation is based on the following assumptions about adversary capabilities: i) the adversary can modify the watermarked text in a computationally efficient manner (e.g., synonym substitution), ii) the adversary uses LLMs less capable than the target LLM, and/or iii) the adversary can train a detector using given watermarked and non-watermarked texts. We argue that these assumptions are realistic and practical. Otherwise, using more capable LLMs could easily generate high-quality, non-watermarked texts to evade detection, while launching computationally expensive attacks would significantly increase the adversary’s cost. Future research directions include exploring alternative threat models to expand the scope of our analysis.
Causal Analysis of Design Choices. While our watermarker taxonomy provides a clear classification framework, we focus on analyzing the holistic design choices of each watermarked (rather than individual design modules) to elucidate the underlying factors driving our experimental observations. This limitation stems from the relatively small number of watermarkers within each taxonomic category, precluding definitive conclusions about the vulnerability of specific taxonomic elements based on current experimental evidence.
A more granular assessment of specific design choices would ideally involve comprehensive ablation studies, systematically modifying individual design elements and comparing their performance against baseline configurations. However, this approach faces significant practical challenges due to the intricate inter-dependencies and tight coupling among watermarker components. To partially address this challenge, in §4.3, we strategically compare two watermarkers that differ only in the design of one specific component (e.g., context dependency), enabling the examination of the effects of specific design variations. We consider a more systematic causal analysis as future research.
Parameter Tuning. Our comparative analysis employs AUC curves based on default parameter settings across watermarkers. For benchmarking purposes, we evaluate True Positive Rate (TPR) at a fixed False Positive Rate (FPR) of 1%, aligning with established practices in comparative studies. This standardized approach enables systematic assessment of watermark detection effectiveness while maintaining consistent false positive control. While this approach provides a pragmatic framework for comparative evaluation, it underscores the importance of future research into comprehensive parameter optimization. Such studies could reveal the full performance potential of each method and yield deeper insights into their relative strengths and limitations.
Ethics Consideration
In this paper, we conduct a systematic study of state-of-the-art LLM watermarkers, focusing on their robustness against watermark removal attacks. We aim to understand how various design choices affect watermarkers’ resilience and identify best practices for operating watermarkers in adversarial environments.
Stakeholder Considerations. The primary stakeholders affected by this research include users and developers of LLM watermarkers, as well as the broader community relying on these techniques. By identifying the strengths/limitations of various watermarkers, our work could influence the perceived reliability and deployment of LLM watermarkers in critical applications. Conversely, exposing these vulnerabilities allows for the improvement of the current techniques, ultimately contributing to more secure and robust systems.
Potential Harms and Benefits. Exposing vulnerabilities in widely adopted security mechanisms can yield contrasting outcomes. Initially, it may erode trust in LLM watermarkers as reliable safeguards, potentially leading to their underuse and leaving systems exposed to unchecked risks. However, by illuminating these weaknesses, our research catalyzes the development of more robust techniques and fosters a nuanced comprehension of their constraints, ultimately fortifying the long-term security ecosystem.
Future Research and Mitigation. Our study has identified several potential countermeasures to address the vulnerabilities we uncovered. These recommendations are designed to steer future research and assist developers in bolstering the security of LLM watermarkers.
@@ Appendix A Parameter Setting
Appendix B Taxonomies
Index-dep. Text-dep. Context-free Dist.-shift Dist.-reweight Dist.-transform Score- Diff.- Model-based
[…]
Table 6: A taxonomy of LLM watermarkers. References: TGRL Kirchenbauer et al. (2023a), UG Zhao et al. (2023), UPV Liu et al. (2023a), SIR Liu et al. (2024), RDF Kuditipudi et al. (2023), UB Hu et al. (2024), DIP Wu et al. (2024), GO Aaronson and Kirchner , CTWL Wang et al. (2024), MPAC Yoo et al. (2024).
B.1 Context Dependency
$r_{t}=f_{k}(t),$ (2)
that only depends on the index $t$ of the next token; $r_{t}$ is then used to generate the perturbation $\Delta_{t}$ (Kuditipudi et al., 2023).
Text-dependent watermark. The watermarker considers the previous tokens $x_{<t}$ as the context window to generate $\Delta_{t}$ (Kirchenbauer et al., 2023a; Aaronson and Kirchner, ; Hu et al., 2024). For instance, GO (Aaronson and Kirchner, ) computes the hash of the concatenation of previous $w$ tokens as $r_{t}$:
$r_{t}=f_{k}(x_{t-w}\|\ldots\|x_{t-1}),$ (3)
while TGRL (Kirchenbauer et al., 2023a) suggests using the min hash of the previous token:
$r_{t}=\min(f_{k}(x_{t-w}),\ldots,f_{k}(x_{t-1}))$ (4)
Similarly, UB (Hu et al., 2024) concatenates the previous $w$ tokens to generate the context code for reweighting the logits. UPV (Liu et al., 2023a) and SIR (Liu et al., 2024) use neural networks to generate $\Delta_{t}$ based on the previous tokens.
Context-free watermark. The watermarker applies a universal perturbation $\Delta_{t}$ across the next-token distributions of all the tokens without considering their contexts (Zhao et al., 2023).
B.2 Generation Strategy
Distribution shift. TGRL (Kirchenbauer et al., 2023a) modifies $p(x_{t}|x_{<t})$ by adding a shift $\delta$ to the logits of “green-list” tokens (the remaining as “red-list” tokens) as the modified distribution $\tilde{p}(x_{t}|x_{<t})$. A token $x$ is considered as green listed if $\pi_{r_{t}}(x)<\gamma d$ where $\pi_{r_{t}}$ is a permutation seeded by $r_{t}$, $\gamma$ is a parameter to control the size of the green list, and $d$ is the number of vocabulary size. Similarly, UG (Zhao et al., 2023) uses a fixed red-green split over the vocabulary, showing greater robustness than TGRL against edit-distance-bounded attacks due to its “hard” split. UPV (Liu et al., 2023a) selects the top-$k$ tokens from $p(x_{t}|x_{<t})$, applies a neural network to predict the green-list tokens from these tokens, and adds $\delta$ to the logits of green-list tokens. Unlike the other strategies, the distribution-shift strategy preserves the diversity of generated tokens; however, it can not be made indistinguishable, as $\tilde{p}(x_{t}|x_{<t})$ and $p(x_{t}|x_{<t})$ are inherently distinguishable.
Distribution reweight. Similar to distribution shift, this strategy alters the next-token distribution but uniquely perturbs the logit of each token. For instance, SIR (Liu et al., 2024) trains a neural network to predict the perturbation to logit of each token, which is unbiased (no preference over specific tokens) and balanced (with total perturbation summing up to 0). UB (Hu et al., 2024) advocates unbiased watermarking such that the expectation of the reweighted distribution agrees with the original distribution. It proposes two reweighting schemes: $\delta$-reweighting uniformly samples a token from $p(x_{t}|x_{<t})$ and changes its probability to 1; $\gamma$-reweighting shuffles all the tokens, rejects the first half, and double the probabilities of the remaining half. DIP (Wu et al., 2024) also uses a distribution-reweight generation strategy similar to UB but does not need to access the LM during detection.
Distribution transform. Another line of watermarkers apply randomized transform on $p(x_{t}|x_{<t})$ to sample $x_{t}$. For instance, RDF-EXP (Kuditipudi et al., 2023) and GO (Aaronson and Kirchner, ) use the Gumbel-max trick and apply the exponential transform. Let $p(x_{t}|x_{<t})=\{p_{i}\}_{i=1}^{d}$ be the distribution over the next token $x_{t}$. Then $x_{t}$ is sampled as:
$x_{t}=\arg\max_{i\leq d}r_{i}^{1/{p_{i}}}$ (5)
where $r_{i}$ is generated by the pseudo-random functions in Eq. 2 and Eq. 3. Similarly, RDF-IST (Kuditipudi et al., 2023) applies inverse transform over $p(x_{t}|x_{<t})$. With $\pi_{k}$ as a random permutation seeded by the secret key $k$, the next token $x_{t}$ is selected as:
B.3 Detection Method
Score-based detection. The detector computes the random value $r_{i}$ for each position, the per-token statistics $s(x_{i},r_{i})$, and a score over such statistics, which is then subjected to a one-tailed statistical test to determine whether the text is watermarked. The per-token statistics vary with the concrete generators. For instance, GO (Aaronson and Kirchner, ) defines $s(x_{i},r_{i})=-\log(1-h_{r_{i}}(x_{i}))$, while TGRL (Kirchenbauer et al., 2023a) defines $s(x_{i},r_{i})=1$ if $x_{i}$ is in the green list and 0 otherwise. One simple way to aggregate the per-token statistics is to compute their sum (Aaronson and Kirchner, ; Kirchenbauer et al., 2023a; Zhao et al., 2023; Liu et al., 2024):
$S=\sum_{i=1}^{n}s(x_{i},r_{i})$ (7)
However, as the random values may be misaligned with the tokens (e.g., due to editing), a more robust way is to compute the alignment score (e.g., edit score (Kuditipudi et al., 2023; Wang et al., 2024)):
$S$ $=s^{\psi}(n,n)$
[…]
where $\psi$ is the “edit-cost” parameter.
Differential-based detection. This line of detectors also relies on the score of a given text. However, the score is computed by comparing the given text with the non-watermarked text generated by the same LLM. For instance, UB (Hu et al., 2024) computes the log-likelihood ratio (LLR) score:
$s(i)=\log\frac{\tilde{p}(x_{i}|x_{<i})}{p(x_{i}|x_{<i})}$ (12)
Model-assisted detection. Instead of computing the per-token statistics, which are often subject to watermark removal attacks, one may also train a model to predict whether the given text is watermarked. UPV (Liu et al., 2023a) trains a neural network, which shares the same embedding layers with its generator, to detect watermarked texts. Similarly, one may develop a generic detector by training it to distinguish watermarked and non-watermarked texts. We explore this option in §4.
Appendix C Fidelity Metrics
In this paper, to evaluate the impact of watermarking on text quality, we employs the following metrics to measure the difference between the original text $T$ and the watermarked text $\tilde{T}$. Below, we provide a detailed description of each metric.
BLEU (Papineni et al., 2002) measures the lexical similarity of $T$ and $\tilde{T}$ by calculating the proportion of $n$-grams matched in $\tilde{T}$ and $T$.
BERTScore (Zhang et al., 2020) measures the token-level similarity of $T$ and $\tilde{T}$ by leveraging the pre-trained contextual embeddings from BERT (Devlin et al., 2019) and matching tokens in $T$ and $\tilde{T}$ using cosine similarity.
P-SP (Wieting et al., 2022) evaluates the semantic similarity of $T$ and $\tilde{T}$ using the cosine similarity of their encodings. In particular, the encoding of $T$ (or $\tilde{T}$) is calculated by averaging the embeddings of its subword units generated by SentencePiece (Kudo, 2018).
MAUVE (Pillutla et al., 2021) compares the distributions of $T$ and $\tilde{T}$ by computing an information divergence curve within a quantized embedding space. The area under this divergence curve provides a scalar summary of the trade-off between Type I ($T$ places high mass in areas where $\tilde{T}$ has low mass) and Type II (vice versa) errors.
Appendix D Details of Watermark Removal Attacks
Linguistic variation attack. This class of attacks perturb the linguistic features of the watermarked text, without changing its semantics. We consider the set of perturbations in HELM (Liang et al., 2023), which simulate natural linguistic variations encountered in human interactions with text typing interfaces: i) Lowercasing converts all the words to lower-cases, which potentially affects the interpretation of proper nouns or emphases. ii) Contracting/expanding replaces phrases with their contracted/expanded forms (e.g., “I am” to “I’m” and vice versa), which may impact the tokenizer.
Lexical editing attack. This class of attacks modifies individual words, aiming to maintain the original text’s semantics. Specifically, we consider the following editing operations: i) Misspelling, similar to text-bugger (Li et al., 2019), replaces words with their common misspellings (plural forms also considered); ii) Typoing replaces certain letters in a word with others; iii) Synonymizing replaces words with their synonyms using WordNet (Miller, 1995); and iv) Swapping randomly exchanges the positions of two words within the text, which alters the text structure while potentially preserving the overall semantics.
Text-mixing attack. This class of attacks aims to “dilute” the watermark by mixing the watermarked text with non-watermarked text fragments. Specifically, the copy-pasting attack (Kirchenbauer et al., 2023b) embeds the watermarked text into the context of non-watermarked, human-written text. Note that the influence of non-watermarked text can be controlled by setting the fractions of watermarked and non-watermarked text fragments (i.e., the mixing weights).
Paraphrasing attack. This class of attacks relies on an additional LLM (i.e., paraphraser) to re-write the given watermarked text to evade the detector. For instance, a light paraphraser (Sankar Sadasivan et al., 2023) (e.g., T5-based paraphraser (Damodaran, 2021)) can paraphrase the watermarked text sentence-by-sentence, while a more capable paraphraser (e.g., DIPPER (Krishna et al., 2023)) paraphrases the watermarked text in one-shot, also enabling to control lexical diversity and token order diversity.
Similar to paraphrasing, the translating attack uses a translator LLM (e.g., Seamless  (Communication et al., 2023)) to cycle the watermarked text through multiple languages (e.g., from English to French and back to English). This process can significantly alter the sentence structure and phrasing.
In this class of attacks, we assume the adversary has access to a generic detector that is trained to distinguish watermarked and non-watermarked texts. The adversary then perturbs the watermarked text based on this surrogate detector. We explore this attack type in §5.
@@ Appendix D Details of Watermark Removal Attacks
Appendix E Watermark Effectiveness
We first evaluate the effectiveness of different watermarkers. Following prior work (Kirchenbauer et al., 2023a; Piet et al., 2023; Kuditipudi et al., 2023), for each watermarker, we sample 1,000 prompts and use the LLM in combination with the watermarker to generate the watermarked texts; meanwhile, we select human responses to the same prompts as the non-watermarked texts. We then measure the accuracy of the watermarker’s detector in distinguishing the watermarked and non-watermarked texts.
E.1 Overall Effectiveness
@@
Meanwhile, most methods observe marginal performance drops on Llama2-7B, as shown in Figure 10 (c-d). For instance, compared with its performance on OPT-1.3B, the AUC of SIR drops by 0.11 and 0.08 on C4 and HC3 respectively. This observation aligns with previous research  (Kuditipudi et al., 2023), indicating that watermarkers are more effective on OPT compared with Llama2. This phenomenon can be partly understood through the following explanation. In contrast of less capable LLMs (e.g., OPT-1.3B), Llama2-7B typically produces texts of lower perplexity. Since most watermarkers inject watermarks by slightly altering the next-token distribution, the lower perplexity in Llama2’s outputs hampers the effectiveness of such perturbations. Moreover, it is observed that the performance of various methods varies significantly across different datasets. For instance, UG’s AUC differs by 0.28 between C4 and HC3, while UB’s AUC differs by 0.11. This observation is further supported by the TPR measures at fixed FPRs (fixed as 1%), as shown in Figure 10 (e-f). Our findings suggest that the concrete dataset/task tends to have a larger impact on watermarkers over more capable LLMs.
E.2 Impact of Text Length
@@
E.3 Impact of Temperature
@@
Figure 12 compares how the TPRs of different watermarkers vary with the setting of $\tau=$ 0.7, 1.0, and 1.3 (with FPRs fixed as 1%). Observe that the performance of most watermarkers marginally improves with $\tau$, which corroborates prior work (Piet et al., 2023). For instance, the TPR of TGRL increases by about 0.05 as $\tau$ varies from 0.7 to 1.3. Interestingly, in contrast, the TPR of UPV decreases as $\tau$ grows. This can be explained as follows. UPV employs a neural network as the detector that depends on general textual features, which tends to be more sensitive to increasing randomness, compared with other watermarkers that rely on specific watermark signals (e.g., green/red-listed tokens).
Appendix F Watermark Fidelity
@@
We evaluate the impact of different watermarkers on the text quality. We compare the original text $T$ and watermarked text $\tilde{T}$ using the metrics (detailed in §3.2) of BERTScore (Zhang et al., 2020), P-SP (Wieting et al., 2022), and MAUVE (Pillutla et al., 2021). The results are summarized in Figure 13.
@@
Appendix G Additional Results
G.1 Multi-bit Watermarking
While our study focuses on one-bit watermarkers, for completeness, we also evaluate CTWL (Wang et al., 2024), a multi-bit watermarker. In contrast of one-bit watermarkers that encode only a single bit of information (i.e., whether a given text is watermarked), a multi-bit watermarker can encode multiple bits of information into the watermarked text, such as the generating model, the date of generation, and other details. However, despite its larger information capacity, we find that a multi-bit watermarker is typically less robust compared to one-bit watermarkers, as illustrated in Figure 14 and  15.
@@ G.1 Multi-bit Watermarking
G.2 Robustness of Watermarkers on Llama2
Our evaluation in §4 mainly uses OPT-1.3B as the underlying LLM. Here, we present the results on Llama2-7B, a more capable LLM. The results can be found at Table  7
Water marker CLEAN Linguistic Variation Textual Integrity Content Manipulation Paraphrase
Contra Expan LowCase Swap Typo Syno Missp CP1-10 CP3-10 CP1-25 CP3-25 DP-20 DP-40 Trans
TGRL 0.896 0.864 0.892 0.75 0.408 0.34 0.818 0.78 0.014 0.074 0.088 0.348 0.536 0.334 0.196
UG 0.814 0.791 0.823 0.697 0.843 0.357 0.75 0.755 0.025 0.054 0.050 0.190 0.771 0.633 0.344
UPV 0.588 0.502 0.598 0.494 0.388 0.018 0.512 0.392 0.005 0.002 0.002 0.007 0.362 0.355 0.042
RDF 0.790 0.735 0.725 0.622 0.333 0.888 0.738 0.824 0.021 0.045 0.045 0.307 0.471 0.303 0.010
UB 0.946 0.926 0.940 0.512 0.016 0.030 0.614 0.650 0.000 0.000 0.000 0.022 0.143 0.036 0.094
SIR 0.566 0.066 0.078 0.078 0.076 0.010 0.064 0.050 0.034 0.026 0.042 0.048 0.040 0.081 0.100
GO 0.996 0.996 0.996 0.959 0.643 0.556 0.989 0.987 0.058 0.228 0.286 0.825 0.900 0.623 0.814
Table 7: Resilience of LLM watermarkers against attacks on Llama2.
G.2.1 Linguistic Variation Attack
@@
G.2.2 Lexical Editing Attack
@@
G.2.3 Text-mixing Attacks
@@
G.2.4 Paraphrasing Attack
@@
G.3 Fidelity Preservation of attacks
@@
Appendix H Additional Related Work
We survey the relevant literature in the following categories: i) detection of LLM-generated texts, ii) LLM watermarking, iii) attacks on LLM watermarking, and iv) evaluation of LLM watermarkers.
Detection of LLM-generated texts. The advances in LLMs also give rise to their possible misuses. There is thus a pressing need for the capability of distinguishing LLM- and human-generated texts. Initial work attempts to either train classifiers using LLM- and human-generated textsMitchell et al. (2023) or to leverage intrinsic characteristics of LLM-generated texts (e.g., perplexity and variability in length, complexity, and information density)Lütkebohle . Yet, with LLMs becoming increasingly capable, the difference between LLM- and human-generated texts is narrowing, making such approaches less effective.
LLM watermarking In response, LLM watermarking emerges as a promising alternative, which instruments the LLM generative process with statistical signals that can be subsequently detected. The existing LLM watermarking techniques can be categorized based on the stages in which they are appliedLiu et al. (2023b): i) training-time watermarkingLiu et al. (2023c); Tang et al. (2023); Sun et al. (2022, 2023); Xu et al. (2024), ii) watermarking during logit-generationKirchenbauer et al. (2023a); Zhao et al. (2023); Hu et al. (2024); Liu et al. (2023a, 2024); Fairoze et al. (2023); Ren et al. (2023); Fernandez et al. (2023); Wang et al. (2024); Yoo et al. (2023); Lee et al. (2023); Giboulot and Teddy (2024), iii) watermarking during token-sampling Kuditipudi et al. (2023); Aaronson and Kirchner ; Hou et al. (2023), and iv) post-generation watermarkingZhang et al. (2023); Yoo et al. (2024); Yang et al. (2023); Munyer and Zhong (2023); Sato et al. (2023); Abdelnabi and Fritz (2021). This study mainly focuses on training-free, pre-generation watermarking, which applies to any given LLMs and provides flexible control over multiple criteria (e.g., quality, effectiveness, and robustness).
Attacks on LLM watermarking. One critical property of an LLM watermarker is its robustness against potential attacks. A variety of attacks can be applied to LLM watermarking, ranging from removing the embedded watermark to uncovering the green/red lists. For instance, DipperKrishna et al. (2023) is a widely used paraphrasing attack to evaluate the robustness of LLM watermarkers Kirchenbauer et al. (2023a); Zhao et al. (2023) against watermark removal attacks; the watermark stealing attackJovanović et al. (2024) is proposed to identify green-list tokens and to replace them with red-list tokens, targeting TGRLKirchenbauer et al. (2023a) and UGZhao et al. (2023) to further launch spoofing or removal attacks. This study primarily focuses on watermark removal attacks as they can target any LLM watermarkers and have profound implications in practice (e.g., disinformation, academic cheating, and automated phishing).
Evaluation of LLM watermarkers. As LLM watermarkers become prevalent, recent work attempts to benchmark the performance of various watermarkers. However, current studies either focus solely on the effectiveness of watermarking or have limited assessments of robustness. For instance, WaterBenchTu et al. (2023) compares the effectiveness of TGRLKirchenbauer et al. (2023a) and UGZhao et al. (2023) under varying hyper-parameter settings (e.g., prompt length); LLM-judgerSingh and Zou (2023) uses GPT-3.5-Turbo as a judge to evaluate the effectiveness of RDFKuditipudi et al. (2023) and TGRLKirchenbauer et al. (2023a) and employs a binary classifier based on MLP to distinguish between watermarked and non-watermarked texts; MarkMyWordsPiet et al. (2023) compares the effectiveness of four watermarkers including TGRLKirchenbauer et al. (2023a), GOAaronson and Kirchner , RDFKuditipudi et al. (2023), and UWChrist et al. (2023) and only evaluates the robustness of TGRL against watermark removal attacks.
Appendix I Evaluation Guidelines for Future LLM Watermarking Research
Robustness. It is often critical to properly set the decision threshold for a watermarker (and also the attacks) to fully assess its robustnessLukas et al. (2021), which unfortunately is often missing in the original papers. To overcome this issue, we recommend i) measuring the overall effectiveness (TPR) in terms of ROC (across different threshold settings), ii) measuring the TPR under a fixed FPR (e.g., 0.01), and iii) considering varying attack intensity.
Fidelity. It is notoriously challenging to meaningfully measure the quality of text dataPillutla et al. (2021). We recommend employing a variety of metrics (e.g., BERTScore, P-SP, MAUVE) to comprehensively measure the quality retention of watermarkers as well as attacks. In addition, one may also leverage external tools (e.g., more advanced LLMs such as GPT-4) or human evaluation to provide a more accurate assessment if feasible.